OpenStudy (anonymous):

What is the constant of variation for the quadratic variation? x 2 3 4 5 6 y 20 45 80 125 180 A. 5 B. 10 C. 15 D. 25

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

A.5

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

I have another

OpenStudy (dumbcow):

y = kx^2 plug in a point 20 = k(2^2) = 4k

OpenStudy (anonymous):

What is the constant of variation for the quadratic variation? 6y = 9x2

OpenStudy (triciaal):

if you have more open a new question

OpenStudy (anonymous):

Oh okay

OpenStudy (luckycoins888):

@k_lynn